What is the theoretical yield of ethyl chloride in the reaction of 19.2 g of ethylene with 52 g of hydrogen chloride? (For ethylene, MW=28.0amu; for hydrogen chloride, MW=36.5amu; for ethyl chloride, MW=64.5amu.)
H2C=CH2+HCl→CH3CH2Cl
1
Expert's answer
2019-10-26T02:58:44-0400

"\frac{(19.2 \ g \ H_2C=CH_2)}{(28.0\ g\ H_2C=CH_2/mol)}""= 0.68571 \ mol \ H_2C=CH_2"

"\frac{(52 \ g\ HCl)}{(36.5 g HCl/mol)}" "= 1.4247 \ mol\ HCl"

"0.68571" mole of "H_2C=CH_2" would react completely with "0.68571" mole of "HCl" , but there is more "HCl" present than that, so "HCl" is in excess and "H_2C=CH_2" is the limiting reactant.

The reaction is :

"H_2C=CH_2+HCl\u2192CH_3CH_2Cl"

Moles of ethene"=" Moles of "CH_3CH_2Cl" formed.

"(0.68571\ mol\ H_2C=CH_2)" "\times" "(64.5 \ g \ CH_3CH_2Cl/mol) ="

"44.2\ g\ CH_3CH_2Cl" in theory.